I can't figure out this problem please explain it to me.

6. A lot is 40 meters by 36 meters. a house 27 meters by 9 meters is built on the lot. How much area is left over for a lawn?
thanks Janelle

Answer by checkley71(8403)

40*36=1440 m^2 is the lot size.
27*9=243 m^2 is the size of the house.
1440-243=1197 m^2 is the size of the area left for a lawn.